# Cultural Transmission, Technology, and Treatment of the Elderly

### v1

@@ 4.1.1 Gifts and transfer of cultural goods
Before continuing to analyze inculcative decisions, we show that the warm glow gift-giving model is equivalent to a model in which cultural goods are transferred between generations. This allows us to relate the importance of gifting in support of the elderly to a more general notion of cultural transmission, its importance in maintaining the elderly, and its importance in society in general.
Suppose that instead of enjoying giving gifts, the current middle-aged generation places value on consumption of a cultural good, for which a price $p_{t}$ must be paid. Suppose that the demand for the cultural good is $x^{d}_{t}$ and that the good is supplied by the current elderly, each of whom has one unit of the good to supply. Lifetime utility (1) in these circumstances could be written as follows:
@@
Given model assumptions, the models captured by utility functions (1) and (8) are observationally equivalent:
  • Gift-giving: Middle-aged agents choose an amount of gifts to give to the elderly, and receive ‘warm-glow’ utility from giving gifts according to the parameter $\eta$.
  • Cultural exchange: Middle-aged agents spend resources $x^{d}_{t}p_{t}$ on cultural goods provided by the elderly, and earn utility from these goods according to the parameter $\eta$.
@@
The gift-giving model resulted in an equilibrium gift given by current middle-aged agents to the elderly that was proportional to current middle-aged disposable income, as in (7). Assuming that inculcation occurs, differentiating (9) with respect to $x^{d}_{t}$ and solving, one finds that the demand for $x^{d}_{t}$ by a current middle-aged is:
@@
By assumption, each elderly agent has a supply of one unit of cultural goods, so demand equal to supply requires $x_{t}^{s}=(1+n)x_{t}^{d}$ or $1=(1+n)x_{t}^{d}$, as each elderly agent possesses a unit of culture. Equating demand from (10) with supply and solving gives $p_{t}=(1+n)\eta(y_{mt}-d_{t})$. From this it follows that $p_{t}x_{t}^{d}=\eta(y_{mt}-d_{t})$ is the payment middle-aged agents make in equilibrium to the elderly for cultural goods, while $p_{t+1}x^{s}_{t+1}=p_{t+1}=(1+n)\eta(y_{mt+1}-d_{t+1})$ is the amount received by the elderly, identical to the gifts implied by $g^{*}_{mt}$ in (7) and the resulting gifts received by the elderly, $g_{et}^{*}$. ∎
The equivalence result in Proposition 2 is useful in that it allows us to work with the easier-to-solve gift-giving model, but also conjecture how the importance of culture interacts with the treatment of the elderly.
@@ 4.1.2 Inculcation decisions
We can also now show that the critical thing for determining inculcation is whether or not elderly income is small relative to middle-aged income. This result and the key income threshold are described in Proposition 2.
@@
Substitution of the optimal gift in (7) back into (5), and substitution of the terminal gift $g^{\prime}_{mt}=\eta y_{mt}$ into (6), coupled with the observation that $g_{et+1}=(1+n)g_{mt+1}$ allows us to write the difference in utilities along the equilibrium path with inculcation, and the utility from discontinuing inculcation, $\Delta U_{t}=U_{mt,\eta}-U_{mt,0}$ as:
@@
Using the assumption that $d_{t}$ is proportional to $y_{mt}$, so $d_{t}=\delta y_{mt}$, where $\delta\in(0,1)$ is a fixed positive constant, and substituting $d_{t}=\delta y_{mt}$ into (12) and rearranging gives:
@@
Proposition 2 defines a critical threshold for the relative income earned by the elderly, so that if relative elderly income falls below the threshold $Y^{*}$, inculcation occurs in equilibrium. The threshold increases (making inculcation more likely) if population growth increases ($n\uparrow$); this is because there are more middle-aged agents for each elderly agent, resulting in more gifts received. Furthermore, if the effectiveness of inculcation increases ($\eta\uparrow$), if inculcation becomes cheaper ($\delta\downarrow$), and if old-age consumption becomes more important ($\beta\uparrow$), inculcation becomes more likely. When inequality (13) does not hold, middle-agers do not anticipate the need to inculcate current youth, as the income they earn when old generates sufficient income for sustenance. At the same time, the relatively low income of middle-agers does not produce sufficiently large gifts to warrant inculcation.
We can now add to the model a measure of the status or prestige of the elderly: the relative consumption of the elderly, $\frac{c_{et}}{c_{mt}}$. Using (7) and (13), we write this consumption ratio as follows, taking into account the range of relative incomes for which inculcation occurs and substituting for equilibrium gifts when necessary:
@@
Equation (14) suggests that there may not be a direct link between inculcative activity and elderly treatment. It may be the case that the elderly provide services or other productive labor of sufficient productive value so that $y_{et}$ is large relative to $y_{mt}$. The elderly will then be well treated, but this could be a result of the nature of production, not of inculcation.
As the ethnographic record emphasizes both the nature of production and different types of property rights as important determinants of the way in which the elderly interact with the economy, particularly among those societies subsisting through hunting and gathering or subsistence agriculture, we now extend the model in this direction.
@@ 4.2 Production Technology
A necessary first step in extending the analysis is the introduction of an aggregate production technology. This allows us to more fully characterize the ultimate sources of income used in expressions such as those in Proposition 2, and how they vary with the base features of the economy; for example, how labor- or land-intensive production is and even be able to distinguish between different types of labor.
Suppose that in each time period, society creates output using labor $L_{t}$ and resources that do not vary with time $T$. We will refer to $T$ as land, but it could be thought of as any other resource in addition to labor that can be owned and used to produce output. Output is produced via the constant-returns production function:
@@
Let $A_{it},i=m,e$ denote the effective labor endowments of middle-aged and elderly agents. $A_{mt}$ might be large relative to $A_{et}$ in situations where most labor is hard physical labor, as is the case among hunter-gatherers who specialize in hunting big game, or in some agricultural societies where farming is difficult and labor intensive work. On the other hand, the elderly may be able to perform some tasks that directly or indirectly provide material gains, such as providing childcare, cooking, or animal processing, which could make $A_{et}$ closer to, but still less than $A_{mt}$. We assume that all agents supply their labor inelastically, so total labor in (15) is effective labor supplies multiplied by population levels:
@@
As the production function (15) is assumed to be constant-returns to scale, it can be decomposed using Euler’s theorem as follows:
@@
The decomposition in (17) gives us a natural way to conceptually attribute output to its sources. We can say $F_{L}L$ is labor’s share of output, while $F_{T}T$ is share created by fixed inputs. Similarly, we can think of the contribution of any agent’s labor to total output as $F_{L}l_{i}$, so we can think of the output created by the labor of different ages of agents as follows:
@@
using this nomenclature, output attributable to the fixed resource is
@@
The role property rights play in the economy is to define how the output created in (18) and (19) is actually allocated to agents, as we now discuss.
@@ 4.3 Property Rights
In Section 3, we emphasized Simmons’ somewhat enigmatic statements about the importance of property rights in providing support for the elderly. Although Simmons suggests that property rights are ‘lifesavers’ for the elderly (Simmons, 1970, p. 36), he also suggests that weaker property rights in some cases are good for the elderly, for example, when food must be shared communally. Our starting point is making a distinction between property rights over inputs such as land and over output - indeed, Simmons seems to confound different conceptions of property rights in his overview of the ethnographic record. Presumably, Simmons meant to say that strong property rights over productive assets are good for the elderly, because asset ownership requires some time and resources be devoted to accumulation or establishment of ownership. But Simmons also seems to indicate that weak property rights over output (i.e., output sharing) also benefit the elderly. This distinction between property rights over output and input, we find, is addressed quite naturally in light of the Euler decomposition of the production function (17).
Let us first imagine that the amount of output created by an agent depends on how much of each productive resource is commanded by the agent. As a starting assumption, suppose that labor is inalienable and therefore any agent is able to command his or her natural share of output $q_{it}^{L}$ from (18). Also, add to this an ownership share of the fixed resource $\sigma_{it},i=m,e$ where $\sum\sigma_{it}=1$ across the population or, in our case, $\sigma_{et}N_{et}+\sigma_{mt}N_{mt}=1$. Then, a starting point to describe the share of output that an agent in some sense creates is expression (20):
$q_{it}=A_{it}F_{L}+\sigma_{it}F_{T}T,\quad i=m,e;\quad\sum_{N}\sigma_{it}=1$ (20)
With (20) as our starting point, we can describe property rights over productive resources using different configurations of $\sigma$, and we can describe property rights over output by connecting $q$, output created, to realized incomes $y_{e}$ and $y_{m}$.
@@ 4.3.1 Property rights over land
The specification in (20) naturally frames our question of the degree to which ownership rights are developed over land (or other fixed resources) in terms of $\sigma$. What values of $\sigma_{it}$ capture the idea that property rights over land are well defined and what values of $\sigma_{it}$ indicate that property rights are not well defined? Consider a situation in which ownership of land is not defined at all, that is, land is a free-access resource. $\sigma_{it}$ should then be set so as to allow our model to mimic a typical common production model, in which agent $i$’s output depends upon $i$’s relative share of labor, so $q_{it}^{L}=\frac{l_{i}}{L}F(L,T)$. [^19] This is a slight generalization of the idea that open access resources generate average product as returns for users as in, for example, Gordon (1954). See Baker and Conning (2024) for further discussion.
Therefore, to reflect the idea that land is completely open access, we use a value of $\sigma_{it}$, $\sigma_{it}^{0}$, as follows:
@@
Where the last part of (22) follows from the assumption that $F(L_{t},T)$ is homogeneous of degree one. This pins one side of a spectrum of possible values for ownership rights over land, with $\sigma_{it}^{0}$ in (21) corresponding to a situation in which there are no land property rights.
[…]
@@
In (21) and (23), we have described free access and fully defined rights over land using values for $\sigma_{it}$ as $\sigma_{it}^{0}$ and $\sigma_{it}^{1}$. We can imagine a spectrum of intermediate possibilities between these two polar extremes and therefore introduce a parameter $\phi_{T}\in[0,1]$ to capture the degree of definition of property rights over land (or other fixed assets). This results in values the following values for $\sigma$:
@@
We can now say when $\phi_{T}=0$ in (24) and (25) there are no ownership rights in land, while when $\phi_{T}=1$ property rights are completely defined, with intermediate values of $\phi_{T}$ corresponding to varying degrees of ownership over land.
@@ 4.3.2 Property rights over output
We have been careful to distinguish the output $q$ actually created by an agent and the income $y$ actually received. We model imperfect rights over output as the loss of some share of $q$ to a general output fund, which is then shared equally among all agents, and do so in such a way so that we can develop expressions for $y$ for each type of agent. Let $\phi$ denote the share of own output retained by the agent so that $1-\phi$ of each agent’s output is contributed to the output pool. Then we can write the total income received by an agent as a function of the degree of definition of property rights as captured by different values of $\phi$ and $\phi_{T}$, assuming that each agent gets an equal share of the total output shared with others:
@@
Substituting into (26) our expressions for incomes in (20) gives:
@@
In (27) we see that imperfect ownership of output removes a bit of output income from the agent but entitles an agent to a share of average output.
@@ 4.3.3 Property rights in general
A final step allows us to discuss the development of property rights in a unified way, rather than discussing property rights in output or inputs. [^21] Developing a theory of the development of property rights is beyond the scope of this paper, but there is no shortage of papers discussing this. For a recent review and one modeling approach, see Baker and Conning (2024). We are supposing that the development of property rights is exogenous in this paper, but typically it is the case that more developed economies have more private property rights. See, e.g., Baker (2008) for a model in which rights over assets emerge endogenously, and Baker and Swope (2021) for a model in which ownership of resources and output-sharing are interrelated. A stylized fact is that property rights tend to develop in tandem; ownership of resources and the output created by those resources seem to evolve in the same direction. We model this fact by assuming that $\phi_{T}$ and $\phi$ are linked. To allow this linkage to be flexible and for the degree of different kinds of property rights to develop at different rates, we introduce $\rho$ as a parameter characterizing the relative speed at which ownership of output and inputs co-evolve and then specify that $\phi_{T}=\phi^{\rho}$.
Since $\phi\in[0,1]$ it follows that when $\rho>1$, $\phi_{T}=\phi^{\rho}<\phi$, and one might say that rights over output are more developed than rights over assets for any interior value of $\phi$. Alternatively, if $\rho<1$, the opposite is true: rights over assets are more developed than rights over output, as $\phi_{T}>\phi$. This approach allows us to acknowledge that different kinds of property rights develop together, while also allowing us to remain agnostic as to which kinds of rights develop faster.
@@ 4.3.4 Back to relative treatment of the elderly
We now can relate the development of property rights to inculcation, treatment of the elderly, and cultural transmission, which means describing how property rights relate to the results in Propositions 1 and 2. Substituting the definitions (24) and (25) into (27), and forming the ratio $\frac{y_{et}}{y_{mt}}$ gives:
@@
Assume that the aggregate production function is Cobb-Douglas: $F(L,T)=L^{\alpha}T^{1-\alpha}$. Then, expression (28) simplifies to:
$\frac{y_{et}}{y_{mt}}=\frac{\phi\left(A_{e}\alpha+\left[A_{e}(1-\phi^{\rho})+\frac{\phi^{\rho}L_{t}}{N_{et}}\right](1-\alpha)\right)+(1-\phi)\frac{L_{t}}{N_{et}+N_{mt}}}{\phi\left(A_{mt}\alpha+A_{mt}(1-\phi^{\rho})(1-\alpha)\right)+(1-\phi)\frac{L_{t}}{N_{et}+N_{mt}}}$ (29)
Expression (29) links relative elderly income, technology of production, and property rights. It provides direct insight into how well the elderly are treated given the fundamental characteristics of the economy on its own, but it also functions as the driving variable in whether or not we expect to see inculcation and a cultural factor in the treatment of the elderly as it appears in the inequality (13) in Proposition 2.
When $\phi=0$, (29) collapses to unity, as one would expect. This is because when all output is shared equally, elderly and middle-aged are the same. In Proposition 3, we collect and further define some of the key results concerning relative income and the definition of property rights.
@@
The relative income of the elderly, $\frac{y_{et}}{y_{m}t}$ is U-shaped in the degree of definition of property rights $\phi$, provided a) production is not too labor-intensive ($\alpha$ is not too large), and provided $A_{e}<A_{m}$; the elderly have a relatively smaller endowment of labor.
@@
As the derivative is initially negative, equals zero at a unique point $\phi^{*}\in(0,1)$ under the maintained assumptions, and then becomes positive, it follows that relative income is U-shaped in the degree to which property rights are developed, $\phi$. ∎
The proposition starts with the assumption that $A_{et}<A_{mt}$, which means that an elderly agent provides fewer effective units of labor than a middle-aged agent. If this assumption did not hold, the relative income of the elderly would be strictly increasing in $\phi$. Moreover, if $\alpha$ were very large, $\phi^{*}$ would be outside the $(0,1)$ range, so relative elderly income only decreases in $\phi$. Note that when $\phi=1$ we have:
@@
The limit ratio in (34) can be greater than one or less than one, depending on the relative elderly population and their relative productivity. For small enough $\alpha$ (that is, production is resource intensive), the ratio in (34) can be greater than one even if $A_{et}=0$ (when the elderly do not supply any material labor to the economy). In these circumstances, the elderly can achieve a high level of prestige measured by relative income even in the absence of inculcation. This is because the elderly have accumulated crucial resources while in middle age.
@@
In Figure 1, the relative income ratio (29) is plotted as a function of the degree of development of property rights, $\phi$, for different assumptions about the underlying nature of production. Although changes in model parameters influence its exact position, it retains the same overall shape. The figure is suggestive of the empirical evidence discussed in the previous section. For one, it shows how, possibly, in a relatively egalitarian society, there is no need for cultural protection of the elderly, but as property rights develop, at least initially, a demand for some other means of ensuring the relative livelihood of the elderly arises as relative consumption of the elderly experiences downward pressure, and developing culture is a way to accomplish this. The provision of cultural goods and things like food taboos is a way of transferring current income for future consumption via inculcation. However, as the possibilities for the ownership of resources become more expansive, this helps the elderly as they are able to benefit from the income that owned assets provide. This may dull the need for direct institutions, such as food taboos, while at the same time the prestige of the elderly may continue to increase.
In Figure 2, we illustrate some possibilities graphically. The figure essentially combines the results of Propositions 2 and 3. As property rights develop ($\phi$ grows larger) and the economy moves to the right along the $x$ axis, $\frac{c_{et}}{c_{mt}}$, which is equivalent to $\frac{y_{et}}{y_{mt}}$ in this range, initially falls and passes the $Y^{*}$ threshold in Proposition (2) at the point marked by the vertical dashed line, so inculcation of young people to give gifts to the elderly begins. This mitigates the decline in the relative treatment of the elderly in the indicated range. However, inculcation ceases once property rights become sufficiently developed, as indicated in the figure.
This analysis suggests that it might be difficult to forge a direct link between culture, inculcation, and the prestige of the elderly, and that any relationship between culture, the means of production, and elderly treatment may be nonlinear.

### v2

@@ 4.1.1 Gifts and transfer of cultural goods
Before continuing to analyze inculcative decisions, we show that the warm glow gift-giving model is equivalent to a model in which cultural goods are transferred between generations. This allows us to relate the importance of gifting in support of the elderly to a more general notion of cultural transmission and its importance in maintaining the elderly.
Suppose that instead of enjoying giving gifts, the current middle-aged generation places value on consumption of a cultural good, for which a price $p_{t}$ must be paid. The demand for cultural goods is $x^{d}_{t}$ and the good is supplied by current elderly, each of whom has one unit of the good to supply. Lifetime utility (1) in these circumstances is:
@@
The models captured by utility functions (1) and (8) are observationally equivalent:
  • Gift-giving: Middle-aged agents choose gifts to the elderly, and receive ‘warm-glow’ utility from giving gifts.
  • Cultural exchange: Middle-aged agents spend $x^{d}_{t}p_{t}$ on cultural goods supplied by the elderly and earn utility from these goods.
@@
The gift-giving model resulted in an equilibrium gift given by current middle-aged agents proportional to current middle-aged disposable income, as in (7). For the cultural goods model, differentiating (9) with respect to $x^{d}_{t}$ and solving gives the demand for $x^{d}_{t}$ as:
@@
Each elderly agent has a supply of one unit of cultural goods, so demand equal to supply requires $x_{t}^{s}=(1+n)x_{t}^{d}$ or $1=(1+n)x_{t}^{d}$. Equating demand from (10) with supply and solving gives $p_{t}=(1+n)\eta(y_{mt}-d_{t})$. It follows that $p_{t}x_{t}^{d}=\eta(y_{mt}-d_{t})$ are the expenditures of middle-aged agents on cultural goods, while $p_{t+1}x^{s}_{t+1}=p_{t+1}=(1+n)\eta(y_{mt+1}-d_{t+1})$ is the amount received by the elderly, identical to the gifts implied by $g^{*}_{mt}$ in (7) and the resulting gifts received by the elderly, $g_{et}^{*}$. ∎
The equivalence result in Proposition 2 is useful in that it allows us to work with the easier-to-solve gift-giving model while also describing how cultural transmission interacts with elderly treatment.
@@ 4.1.2 Inculcation decisions
We can now show that the inculcation decision depends on whether or not elderly income is small relative to middle-aged income. This result and the key income threshold are described in Proposition 2.
@@
Substitution of the gift in (7) into (5), and substitution of the terminal gift $g^{\prime}_{mt}=\eta y_{mt}$ into (6) coupled with the observation that $g_{et+1}=(1+n)g_{mt+1}$ allows writing the difference in utilities along the equilibrium path with inculcation and the utility from discontinuing inculcation, $\Delta U_{t}=U_{mt,\eta}-U_{mt,0}$ as:
@@
Using the assumption that $d_{t}$ is proportional to $y_{mt}$, so $d_{t}=\delta y_{mt}$, where $\delta\in(0,1)$ is a fixed positive constant, and substituting into (12) gives:
@@
Proposition 2 defines a critical threshold for the relative income earned by the elderly. If relative elderly income falls below the threshold $Y^{*}$, inculcation occurs in equilibrium. The threshold increases (making inculcation more likely) if population growth increases ($n\uparrow$); this is because there are more middle-aged agents for each elderly agent, resulting in more gifts received. If the effectiveness of inculcation increases ($\eta\uparrow$), if inculcation becomes cheaper ($\delta\downarrow$), and if old-age consumption becomes more important ($\beta\uparrow$), inculcation becomes more likely. When inequality (13) does not hold, middle-agers do not anticipate the need to inculcate current youth, as the income they earn when old generates sufficient income for sustenance. Also, the relatively low income of middle-agers does not produce sufficiently large gifts to warrant inculcation.
We now add to the model a measure of the status or prestige of the elderly: the relative consumption of the elderly, $\frac{c_{et}}{c_{mt}}$. Using (7) and (13), we write this consumption ratio as follows, accounting for the range of relative incomes for which inculcation and concurrent gift-giving occurs:
@@
Equation (14) suggests that there may not be a direct link between inculcative activity and elderly treatment. The elderly may provide services or other productive labor so that $y_{et}$ is large relative to $y_{mt}$. The elderly will then be well treated, but this could be a result of the nature of production, not of inculcation.
The ethnographic record highlights production and property rights as key factors shaping the elderly’s economic roles, prompting us to extend the model accordingly.
@@ 4.2 Production Technology
Suppose that in each time period, society creates output using labor $L_{t}$ and resources $T$ that do not vary with time. We will refer to $T$ as land, but it could represent some other resource used to produce output. Output is produced via the constant-returns production function:
@@
Let $A_{it},i=m,e$ denote the effective labor endowments of middle-aged and elderly agents. Agents supply their labor inelastically, so total labor in (15) is effective labor supplies multiplied by population levels:
@@
The production function (15) is constant-returns so it can be decomposed using Euler’s theorem. Using $F_{L}=\frac{\partial F}{\partial L}$ and $F_{T}=\frac{\partial F}{\partial T}$, we may write:
@@
The decomposition in (17) gives us a natural way to attribute output to factors of production; $F_{L}L$ is labor’s contribution to output, while $F_{T}T$ is output created by fixed inputs. Similarly, the contribution of any agent’s labor to output is $F_{L}l_{i}$, so output created by the labor of different ages of agents is:
@@
output attributable to the fixed resource is
@@
Property rights define how the output created in (18) and (19) is allocated to agents.
@@ 4.3 Property Rights
In Section 3, we discussed Simmons’ intriguing remarks on the role of property rights in supporting the elderly. While he describes property rights as ‘lifesavers’ for the elderly (Simmons, 1970, p. 36), he also notes that weaker property rights can sometimes benefit them, such as when food is shared communally. To clarify, we distinguish between property rights over inputs like land and property rights over output such as game acquired. Simmons appears to blur these distinctions in his ethnographic overview. Strong property rights over productive assets likely benefit the elderly by ensuring resources are invested in ownership or accumulation. Conversely, weak property rights over output, encouraging communal sharing, also seem advantageous for them. This input-output distinction aligns naturally with the Euler decomposition of the production function (17).
Suppose that each agent has an ownership share of the fixed resource $\sigma_{it},i=m,e$ where $\sum\sigma_{it}=1$ across the population or, in our case, $\sigma_{et}N_{et}+\sigma_{mt}N_{mt}=1$. Then, a starting point to describe the output that an agent is responsible for creating is expression (20):
$q_{it}=q_{it}^{L}+q_{it}^{T}=A_{it}F_{L}+\sigma_{it}F_{T}T,\quad i=m,e;\quad\sum_{N}\sigma_{it}=1$ (20)
@@ 4.3.1 Property rights over land
The specification in (20) naturally frames property rights over land (or other assets) in terms of $\sigma$. Consider a situation in which ownership of land is not defined at all, that is, land is a free-access resource. $\sigma_{it}$ should then be set so as to allow our model to mimic a typical common production model, in which agent $i$’s output depends upon $i$’s relative share of labor, so $q_{it}=q_{it}^{L}=\frac{l_{i}}{L}F(L,T)$. [^11] This is a slight generalization of the idea that open access resources generate average product as returns for users as in, for example, Gordon (1954). See Baker and Conning (2024) for further discussion.
Open-access land rights is then captured by $\sigma_{it}$ as follows:
@@
Where the last part of (22) follows from the assumption that $F(L_{t},T)$ is homogeneous of degree one. This pins down one side of a spectrum of possible degrees of ownership rights over land, with $\sigma_{it}^{0}$ in (21) corresponding to a situation in which there are no land property rights.
When property rights are fully defined over land, suppose that a right of first possession determines ownership. This can be approximated by the elderly controlling assets in a way proportional to past usage (that is, their usage during middle age), so that when asset rights are fully defined, $\sigma_{et}=\frac{A_{mt-1}}{A_{mt-1}N_{mt-1}}=\frac{1}{N_{et}}$. It follows that in this case the middle-aged do not have any property rights, so $\sigma_{mt}=0$. Fully defined property rights over land are then characterized by:
@@
In (21) and (23), we have described free access and fully defined rights over land using values for $\sigma_{it}$ as $\sigma_{it}^{0}$ and $\sigma_{it}^{1}$. A spectrum of intermediate possibilities exists between the two endpoint cases, and we introduce a parameter $\phi_{T}\in[0,1]$ to capture the degree of definition of property rights over land (or other fixed assets). This results in values the following values for $\sigma$:
@@
When $\phi_{T}=0$ in (24) and (25) there are no ownership rights in land, while when $\phi_{T}=1$ property rights are completely defined.
@@ 4.3.2 Property rights over output
We model imperfect rights over output as the loss of some share of $q$ to a general output fund, which is then shared equally among all agents. Let $\phi$ denote the share of own output retained by the agent so that $1-\phi$ of each agent’s output is contributed to the output pool. Then we can write the total income received by an agent as a function of the degree of definition of property rights as captured by different values of $\phi$ and $\phi_{T}$, assuming that each agent gets an equal share of the total output shared with others:
@@
Substituting into (26) our expressions for outputs in (20) gives:
@@
In (27) we see that imperfect ownership of output removes a bit of output from each agent but entitles an agent to a share of average output.
@@ 4.3.3 Property rights in general
A stylized fact is that property rights tend to develop in tandem; ownership of resources and the output created by those resources seem to evolve in the same direction. [^12] Developing a theory of the development of property rights is beyond the scope of this paper. For a recent review and one modeling approach, see Baker and Conning (2024). In particular see Baker (2008) for a model in which rights over assets emerge endogenously, and Baker and Swope (2021) for a model in which ownership of resources and output-sharing are interrelated. We model this fact by assuming that $\phi_{T}$ and $\phi$ are linked. To allow for different kinds of property rights to develop at different rates yet still be linked, we introduce $\rho$ as a parameter characterizing the relative speed at which ownership of output and inputs co-evolve and then specify that $\phi_{T}=\phi^{\rho}$. When $\rho>1$, $\phi_{T}=\phi^{\rho}<\phi$, and rights over output are more developed than rights over assets for any interior value of $\phi$. Alternatively, if $\rho<1$, the opposite is true: rights over assets are more developed than rights over output, as $\phi_{T}>\phi$.
@@ 4.3.4 Back to relative treatment of the elderly
We now can relate the development of property rights to inculcation, treatment of the elderly, and cultural transmission. Substituting (24) and (25) into (27), and forming the ratio $\frac{y_{et}}{y_{mt}}$ gives:
@@
Assume that the aggregate production function is Cobb-Douglas: $F(L,T)=L^{\alpha}T^{1-\alpha}$. Then, expression (28) becomes:
$\frac{y_{et}}{y_{mt}}=\frac{\phi\left(A_{et}\alpha+\left[A_{et}(1-\phi^{\rho})+\frac{\phi^{\rho}L_{t}}{N_{et}}\right](1-\alpha)\right)+(1-\phi)\frac{L_{t}}{N_{et}+N_{mt}}}{\phi\left(A_{mt}\alpha+A_{mt}(1-\phi^{\rho})(1-\alpha)\right)+(1-\phi)\frac{L_{t}}{N_{et}+N_{mt}}}$ (29)
Expression (29) links relative elderly income, technology of production, and property rights. It provides direct insight into how well the elderly are treated given the fundamental characteristics of the economy on its own, but it also functions as the driving variable in whether or not we expect to see inculcation and a cultural factor in the treatment of the elderly as it appears in the inequality (13) in Proposition 2. When $\phi=0$, (29) collapses to unity. This is because when all output is shared equally, elderly and middle-aged are the same. In Proposition 3, we collect the key results concerning relative income and the definition of property rights embodied in (29).
@@
The relative income of the elderly, $\frac{y_{et}}{y_{m}t}$ is U-shaped in the degree of definition of property rights $\phi$, provided a) production is not too labor-intensive ($\alpha$ is not too large), and b) provided $A_{e}<A_{m}$; the elderly have a relatively smaller endowment of labor.
@@
As the derivative in equation (31) is initially negative, equals zero at a unique point $\phi^{*}\in(0,1)$ and then becomes positive, relative income is U-shaped in the degree to which property rights are developed, $\phi$. ∎
Proposition 3 starts with the assumption that $A_{et}<A_{mt}$, which means that an elderly agent provides fewer effective units of labor than a middle-aged agent. Another assumption is that $\alpha$ is not very large; if it were $\phi^{*}$ would be outside the $(0,1)$ range and relative elderly income would only decreases in $\phi$. When $\phi=1$ we have:
@@
The limit ratio in (34) can be greater than one or less than one, depending on the relative elderly population and their relative productivity. For small enough $\alpha$ (that is, production is resource intensive), the ratio in (34) can be greater than one even if $A_{et}=0$ (the elderly do not supply any material labor). Then, the elderly can achieve a high level of prestige measured by relative income even in the absence of inculcation. This is because the elderly have accumulated assets in middle age.
@@
Figure 1 plots the relative income ratio (29) as a function of property rights development, $\phi$, under varying assumptions about production. While model parameters can shift the curve’s position, its overall shape remains consistent. The figure aligns with the empirical evidence discussed earlier. It suggests that in relatively egalitarian societies, cultural protections for the elderly may be unnecessary. However, as property rights initially develop, downward pressure on the relative consumption of the elderly creates demand for mechanisms like cultural inculcation to ensure their livelihood. For example, cultural goods or food taboos can transfer current income into future consumption. As resource ownership expands, asset income begins to benefit the elderly directly, reducing the need for such institutions, even as their prestige continues to rise.
Figure 2 combines results from Propositions 2 and 3. As property rights develop ($\phi$ increases) and the economy moves rightward along the $x$-axis, the relative consumption ratio $\frac{c_{et}}{c_{mt}}$ (equivalent to $\frac{y_{et}}{y_{mt}}$ in this range) initially declines, crossing the $Y^{*}$ threshold from Proposition 2 at the vertical dashed line. At this point, inculcation begins, encouraging young people to support the elderly, mitigating the decline in relative treatment. However, once property rights become sufficiently advanced, inculcation ceases, as shown in the figure. This analysis highlights the complexity of linking culture, inculcation, and elderly prestige, suggesting a nonlinear relationship between culture, production, and elderly treatment.